# Casper: Causality-Aware Spatiotemporal Graph Neural Networks for Spatiotemporal Time Series Imputation

### v1

@@ 4.4. Visualization
4.5. Sensitivity Analysis
We present the results of sensitivity experiments of $\lambda$ and the number of prompts $N_{P}$ in Figure 6(a)-6(b). For $\lambda$, the lowest MAE can be obtained when $\lambda$ is around 0.001. For $N_{P}$, the best results can be obtained when $N_{P}\in[200,1400]$.
@@ 5. Related Work
[…]
@@
Causal Inference. Causality theory (Pearl and Mackenzie, 2018) provides theoretical guidance to design causality aware models. It has been widely explored in the computer vision domain to discover causal relationships (Bengio et al., 2019), generate counter-factual samples (Abbasnejad et al., 2020; Yue et al., 2021; Kocaoglu et al., 2018) and reduce bias (Hu et al., 2021; Yang et al., 2021; Qi et al., 2020). In the graph mining domain, CGI (Feng et al., 2021) studies how to select trustworthy neighbors during inference; CLEAR (Ma et al., 2022a) explores how to generate counterfactual explanations for graph-level prediction models based on Independent Component Analysis (ICA) (Khemakhem et al., 2020); HyperSCI (Ma et al., 2022b) explores Individual Treatment Effect (ITE) on hyper-graphs. NEAT (Ma et al., 2023) investigates the impact of MRSA infection via the Neyman-Rubin potential outcome framework (Rubin, 2005). CAL (Sui et al., 2022) introduces a causal attention learning framework for graph neural networks based on the backdoor adjustment (Pearl and Mackenzie, 2018). There are two differences between Casper and the above graph methods: (1) our setting is the dynamic spatio-temporal setting but their setting is static graph; (2) Casper is based on the frontdoor adjustment and Granger causality, which is fundamentally different from their theoretical basis of causality. In the time series domain, the Granger causality (Granger, 1969) is widely used for analyzing the causality between time series in the forecasting setting. GrID-Net (Wu et al., 2021) leverages the Granger causality to infer regulatory locus–gene links. cLSTM (Tank et al., 2021) and economy-SRU (Khanna and Tan, 2019) integrates the Granger causality with LSTM (Hochreiter and Schmidhuber, 1997) and SRU (Oliva et al., 2017). However, thses methods require the input time series to be fully observed. CUTS (Cheng et al., 2022) is a recently proposed two-stage model, which first imputes missing data and then discovers causality between time series. There are three major differences between Casper and CUTS. (1) CUTS does not consider confounders, while Casper removes confounders via the frontdoor adjustment. (2) CUTS has to re-train the causal model for each input snippet, but Casper does not have such a requirement. (3) CUTS is a two stage model, while Casper is an end-to-end model.
@@ 6. Conclusion
In this paper, we review the spatiotemporal time series imputation task via the Structure Causal Model (SCM), which shows the causal relationships among the input, output, embeddings and confounders. The confounders could open shortcut backdoor paths between the input and output, which could mislead the model to learn the non-causal relationships. We use the frontdoor adjustment to block the backdoor paths. Based on the results of the frontdoor adjustment, we proposed a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper), which is comprised of a Prompt Based Decoder (PBD) and an encoder equipped with Spatiotemporal Causal Attention (SCA). The proposed PBD could reduce the impact of the confounders at a high level. The proposed SCA could further discover the sparse causal relationships among embeddings. Theoretical analysis shows that SCA decides causal and non-casual relationships based on the values of gradients. Experimental results on three real-world benchmark datasets show that Casper could outperform the baselines methods, and could also effectively discover the causal relationships.

### v3

@@ 5. Related Work
In this section, we briefly review the most relevant works to ours, including spatiotemporal time series methods as well as causal inference methods.
5.1. Spatiotemporal Time Series Imputation
Spatiotemporal time series imputation is one of the fundamental tasks for time series analysis (Esling and Agon, 2012; Wang et al., 2024; Zheng et al., 2024b). Traditional machine learning approaches are based on statistical analysis, such as linear autoregression (Yi et al., 2016; Durbin and Koopman, 2012) and matrix factorization (Yu et al., 2016). At present, deep learning methods have become the mainstream. Most existing deep learning methods are based on Recurrent Neural Networks (RNN). GRU-D (Che et al., 2018) is one of the first RNN-based imputation models. BRITS (Cao et al., 2018) leverage bi-directional RNN to impute missing data. GAIN (Luo et al., 2018) and E2GAN (Luo et al., 2019) further apply Generative Adversarial Network (GAN) (Goodfellow et al., 2014) to enhance the performance. mTAND (Shukla and Marlin, 2020) adds attention mechanism to RNN. These methods suffer from error propagation and accumulation brought by the auto-regression nature of RNN. To address this issue, non-autoregressive methods are proposed such as NAOMI (Liu et al., 2019), NRTSI (Shan et al., 2023) and the recent Transformer (Vaswani et al., 2017) based methods (Yıldız et al., 2022). There are also some other types of methods, such as Ordinary Differential Equations (ODEs) methods (Rubanova et al., 2019, 2019), state space models (Alcaraz and Strodthoff, 2022) and diffusion models (Tashiro et al., 2021) The above methods mainly focus on the temporal patterns of time series, yet largely ignore the spatial relationships, e.g., distance, among time series. To capture spatial relationships, graph neural networks (Kipf and Welling, 2016; Veličković et al., 2017) are extended to the spatiotemporal setting. LG-ODE (Huang et al., 2020) combines graph neural networks with ODE methods (Rubanova et al., 2019). RETIME (Jing et al., 2022b) introduces a retrieval-based time series model, which leverages retrieved time series as an augmentation for the target time series. NET3 (Jing et al., 2021b) introduces a tensor graph neural network to model the high-order relationships among time series. GRIN (Cini et al., 2021) introduces a bidirectional message passing RNN with a spatial decoder. SPIN (Marisca et al., 2022) presents a sparse spatiotemporal graph neural network for spatiotemporal time series imputation. Recently, PoGeVon (Wang et al., 2023) proposes a position-aware graph neural network based variational auto-encoder to impute both time series and edges. However, these methods try to exploit all the available related information for the target missing point, without distinguishing the causal and non-causal relationships, which might have the overfitting problem and make the model vulnerable to noise. Our proposed Casper could distinguish causal and non-causal relationships.
5.2. Causal Inference
@@
Causality theory (Pearl and Mackenzie, 2018) provides theoretical guidance to design causality-aware models. It has been widely explored in the computer vision domain to discover causal relationships (Bengio et al., 2019), generate counter-factual samples (Abbasnejad et al., 2020; Yue et al., 2021; Kocaoglu et al., 2018) and reduce bias (Hu et al., 2021; Yang et al., 2021; Qi et al., 2020). In the graph mining domain, CGI (Feng et al., 2021) studies how to select trustworthy neighbors during inference; CLEAR (Ma et al., 2022a) explores how to generate counterfactual explanations for graph-level prediction models based on Independent Component Analysis (ICA) (Khemakhem et al., 2020); HyperSCI (Ma et al., 2022b) explores the Individual Treatment Effect (ITE) on hyper-graphs. NEAT (Ma et al., 2023) investigates the impact of MRSA infection via the Neyman-Rubin potential outcome framework (Rubin, 2005). CAL (Sui et al., 2022) introduces a causal attention learning framework for graph neural networks based on the backdoor adjustment (Pearl and Mackenzie, 2018). There are two differences between Casper and the above graph methods: (1) our setting is the dynamic spatiotemporal setting but their setting is static graph; (2) Casper is based on the frontdoor adjustment and Granger causality, which is fundamentally different from their theoretical basis of causality. In the time series domain, the Granger causality (Granger, 1969) is widely used for analyzing the causality between time series in the forecasting setting. GrID-Net (Wu et al., 2021) leverages the Granger causality to infer regulatory locus–gene links. cLSTM (Tank et al., 2021) and economy-SRU (Khanna and Tan, 2019) integrates the Granger causality with LSTM (Hochreiter and Schmidhuber, 1997) and SRU (Oliva et al., 2017). However, these methods require the input time series to be fully observed. CUTS (Cheng et al., 2022) is a recently proposed two-stage model, which first imputes missing data and then discovers causality between time series. There are three major differences between Casper and CUTS. (1) CUTS does not consider confounders, while Casper removes confounders via the frontdoor adjustment. (2) CUTS has to re-train the causal model for each input segment, but Casper does not have such a requirement. (3) CUTS is a two-stage model, while Casper is a one-stage end-to-end model.
@@ 6. Conclusion
In this paper, we review the spatiotemporal time series imputation task via the Structure Causal Model (SCM), which shows the causal relationships among the input, output, embeddings, and confounders. The confounders could open shortcut backdoor paths between the input and output, which could mislead the model to learn the non-causal relationships. We use the frontdoor adjustment to block the backdoor paths. Based on the results of the frontdoor adjustment, we propose a novel Causality-Aware Spatiotemporal Graph Neural Network (Casper), which is comprised of a Prompt Based Decoder (PBD) and an encoder equipped with Spatiotemporal Causal Attention (SCA). The proposed PBD could reduce the impact of the confounders at a high level. For SCA, we first extend the definition of Granger causality for time series to embeddings. Then we introduce the architecture of SCA based on the definition, which could discover the sparse causal relationships among embeddings. Theoretical analysis shows that SCA decides causal and non-casual relationships based on the values of gradients. Experimental results on three real-world benchmark datasets show that Casper could outperform the baseline methods for the imputation task. Further analysis shows that Casper could effectively discover the sparse causal relationships.